# Entanglement entropy bounds for pure states of rapid decorrelation

### v1

@@ Appendix A Fannes’ type bound
The control of entropy differences partially relied on the minor modification (4.6) of Fannes’ well-known entropy bound [28] on the difference of von Neumann entropies of two states $\varrho,\widehat{\varrho}$ in terms of their trace distance and the rank of their difference. Trivially estimating $\mathrm{rank}(\varrho-\widehat{\varrho})$ by the dimension of the Hilbert space on which the states are defined, the following Proposition reduces to [28]. Similarly as the latter and unlike the optimal bound by Audenaert [11], the following bound (which repeats (4.6)) is not sharp.
@@
For the reader’s convenience, we include a short proof, which is slightly different than the one in  [28] and which we base on the subadditivity of the trace of the following function of non-negative operators. We set $F:[0,\infty)\to[0,\infty)$ with

### v2

@@ Appendix A Fannes’ type bound
In controling of entropy differences we made use of the following minor modification of Fannes’ entropy bound [29], which appeared above as (4.3), on the difference of von Neumann entropies of two states $\varrho,\widehat{\varrho}$. For the reader’s convenience we include a short proof, which is slightly different than the one in [29].
@@
Bounding $\mathrm{rank}(\varrho-\widehat{\varrho})$ by the dimension of the Hilbert space on which the states are defined, (A.1) reduces to the Fannes bound of [29]. Unlike the optimal bound by Audenaert [11], this bound is not sharp.
Our proof is based on the subadditivity of the trace of the extension of the following function to non-negative operators.
@@
[…]
Declarations The authors declare that they have no competing financial interests or personal relationships that could have appeared to influence the work reported in this paper. Data sharing does not apply to this article as no datasets were generated or analyzed during the current study.